OpenStudy (anonymous):

Which is a solution to the inequality? y < -2/3x + 2 (0, 2) (3, 0) (0, 0) (0, 3)

OpenStudy (misty1212):

hi we can try them !

OpenStudy (misty1212):

if you pput \(x=0,y=2\) you get \[2<0+2\] which is not right so it is not that one (unless it was really \(\leq \))

OpenStudy (misty1212):

if you put \(x=3\) and \(y=0\) up get \[0<-\frac{2}{3}\times 3+2\\ 0<-2+2\\ 0<0\] also not right

OpenStudy (anonymous):

OH i get it now! ill try the last two options myself will you tell me if i get it right?

OpenStudy (anonymous):

sorry i had to log off earlier, but its C right?

sammixboo (sammixboo):

So, is the inequality \(\le\) or \(<\)? Just curious

OpenStudy (anonymous):

the second option :)

sammixboo (sammixboo):

Because sometimes it makes it look weird to make it look like a negative sign

sammixboo (sammixboo):

KK I was curious :) Anyways, you think it is C?

OpenStudy (anonymous):

yes i do think so :0

sammixboo (sammixboo):

You are correct! Plug in the answers and check if you want! :)

OpenStudy (anonymous):

yes thats what i did !! :)

sammixboo (sammixboo):

Okie Dokie! :)

OpenStudy (anonymous):

thank you though btw

sammixboo (sammixboo):

No problem!!

sammixboo (sammixboo):

If you need anymore help tag me @sammixboo or Misty @misty1212